OpenStudy (anonymous):

Differentiate: ln√(9-t^2 ) Differentiate: ln√(9-t^2 ) @Mathematics

OpenStudy (anonymous):

So by the chain rule would i firs differentiate 9-t^2?

OpenStudy (anonymous):

Then differentiate \[(D _{x}[9-t ^{2}])^{1/2}\]

OpenStudy (anonymous):

Then differentiate: ln(Dx[(Dx[9−t2])1/2])

OpenStudy (anonymous):

I would first rewrite the problem. \[\ln((9-t^{2})^{1/2})\] Then using rule of logs, you can take down the exponent to make. \[1/2 \ln(9-t^{2})\] Then we differentiate. When taking derivative of ln, it's just 1/x, rule of logs. So... \[(1/2) * (1/(9-t^{2}) * (2t)\] *remember to use the chain rule when taking the derivative of ln(9-t^2). So clean it up and you have... \[2t/(18-2t^{2})\] ---- I may be wrong. This is what I got though....

OpenStudy (anonymous):

Ah ok

OpenStudy (anonymous):

That rule of logs loves to escape my brain

OpenStudy (anonymous):

Factor out 2 on the bottom and we have t/(9-t^)

OpenStudy (anonymous):

Also, another thing I always forget. Remember the derivative of e^x is e^x. So important to remember.

OpenStudy (amistre64):

differenetiate means to go down as opposed to integrate right?

OpenStudy (amistre64):

\[[ln√(9-t^2 )]'=\frac{[\sqrt(9-t^2)]'}{\sqrt(9-t^2)}\]is a nasic start to it

OpenStudy (amistre64):

*basic start to it

OpenStudy (amistre64):

\[[\sqrt(9-t^2)]'=\frac{[9-t^2]'}{2\sqrt(9-t^2)}\] is anohter part to it

OpenStudy (anonymous):

By differentiate I mean take the derivative of, thats what it means no?

OpenStudy (amistre64):

\[[9-t^2]'=-2t\]is the last bit

OpenStudy (amistre64):

i get so many terms in my head that i get them mixed up alot; they need less d words and more j words perhaps lol

OpenStudy (anonymous):

Hmm, so t/(9-t^) isnt the solution?

OpenStudy (anonymous):

Cause wolfram agrees with t/(9-t^): http://www.wolframalpha.com/input/?i=differentiate%3A+ln%E2%88%9A%289-t%5E2+%29

OpenStudy (amistre64):

dunno, I havent put it all together yet and simplified

OpenStudy (anonymous):

So where did I mess up? That's what I want to know. As for what amistre is doing, i have no idea? Can you explain what you are doing? How did you do your first step?

OpenStudy (amistre64):

\[\frac{[\sqrt(9-t^2)]'}{\sqrt(9-t^2)}\] \[\frac{\frac{[9-t^2]'}{2\sqrt(9-t^2)}}{\sqrt(9-t^2)}\] \[\frac{[9-t^2]'}{{2\sqrt(9-t^2)}*\sqrt(9-t^2)}\] \[\frac{[9-t^2]'}{2(9-t^2)}\] \[\frac{-2t}{2(9-t^2)}\] \[\frac{-t}{9-t^2}\]

OpenStudy (amistre64):

the first step is just to diff the ln|u| into its basic part: du/u

OpenStudy (amistre64):

then its just a matter of following the chain rule for the successive diffs

OpenStudy (anonymous):

hoodrych's methodology seems a bit simpler

OpenStudy (amistre64):

(1/2) * (1/(9-t^{2}) * (2t) <-- should be -2t

OpenStudy (anonymous):

[(√9−t^2)]′ / (√9−t^2 If you're just rewriting du/u then where is the ln on the numerator?

OpenStudy (anonymous):

Good catch amistre.

OpenStudy (amistre64):

there doesnt need to be an ln in the numerator; im just following the ln' rule

OpenStudy (anonymous):

The ln` rule is d/dx ln(x) = 1/x That's not what you did? I'm not saying you're wrong, just trying to understand.

OpenStudy (amistre64):

\[[\ ln|f(u)|\ ]'=\frac{f'(u)}{f(u)}\]

OpenStudy (amistre64):

they make alot of different formats for the same basic rule, if you always remember to just chain out your argument, youll do fine

OpenStudy (anonymous):

So the f`(u)/f(u) is the same thing as the rule I said? (the 1/x)???

OpenStudy (amistre64):

\[\frac{d}{dx}ln(x)=\frac{dx/dx}{x}=\frac{1}{x}\]

OpenStudy (amistre64):

your simply used to ignoring the dx/dx since its equal to 1

OpenStudy (amistre64):

yes, its the same thing you did :)

OpenStudy (anonymous):

aaaaaaaaaaand, click * Thanks bud.

OpenStudy (amistre64):

youre welcome ;)